# The Mittag-Leffler theorem for proper minimal surfaces and directed meromorphic curves

### v1

@@ 2.  Notation and preliminaries
We write =$\mathbb{N}=\{1,2,\ldots\}$ and +=$\mathbb{Z}_{+}=\mathbb{N}\cup\{0\}$. Let ||$\cdot$|| ⋅ | denote the Euclidean norm in $\mathbb{R}^{n}$ for $n\in\mathbb{N}$. Given a compact topological space $K$ and a continuous map $f:K\to\mathbb{R}^{n}$, we mean by $\|f\|_{K}$ the maximum norm of $f$ on $K$. Given $f,g\colon K\to\mathbb{R}^{n}$, we say that $f$ approximates $g$ on $K$, and write $f\approx g$, if $||f-g||_{K}<\varepsilon$ for some $\varepsilon>0$ which is sufficiently small according to the argument. Throughout the paper, we assume that surfaces are connected and $n\in\mathbb{N}$, $n\geq 3$.
Let $A\subset\mathbb{C}^{n}$ be a closed irreducible conical complex subvariety of $\mathbb{C}^{n}$ which is not contained in any hyperplane and is smooth away from 00.333A complex subvariety $A\subset\mathbb{C}^{n}$ is conical if $\lambda A=A$ for all $\lambda\in\mathbb{C}\setminus\{0\}$. We denote by $A_{*}:=A\setminus\{0\}$ the punctured subvariety, assuming that it is smooth and connected.
@@
Let $F=(F_{1},\dots,F_{n})\colon M\to\mathbb{C}^{n}$ be a meromorphic map which is holomorphic on $M\setminus E$ and has effective poles at all points in $E$. Fix a nowhere vanishing holomorphic $1$-form $\theta$ on $M$ (see [18]) and write $dF_{j}=f_{j}\theta$, $j=1,\dots,n$. Then the map $F$ is a meromorphic $A$-immersion $M\setminus E\to\mathbb{C}^{n}$ if and only if its derivative $f=dF/\theta=(dF_{1}/\theta,\dots,dF_{n}/\theta):M\setminus E\to\mathbb{C}^{n}$ has the range in $A_{*}$. On the other hand, given a meromorphic map $f=(f_{1},\dots,f_{n})\colon M\to A_{*}$ with effective poles in $E$, the vectorial $1$-form $f\theta=(f_{1}\theta,\dots,f_{n}\theta)$ integrates to a meromorphic $A$-immersion $F:M\to\mathbb{C}^{n}$ in $\mathscr{I}_{A}(M|E)$ if and only if $f\theta$ is holomorphic and exact on $M\setminus E$. The corresponding meromorphic $A$-immersion is then obtained by
@@
Let $A$ be as in Definition 2.1, $M$ be either an open Riemann surface or a compact bordered Riemann surface, and $\theta$ be a nowhere vanishing holomorphic $1$-form on $M$. A holomorphic $A$-immersion $F\colon M\to\mathbb{C}^{n}$ is full if the $\mathbb{C}$-linear span of the tangent spaces $T_{f(x)}A,\ x\in M$, equals $\mathbb{C}^{n}$. (Here, $dF=f\theta$.) Full meromorphic $A$-immersions are defined analogously.
Given complex manifolds $X,\ Y$ and a subset $M\subset X$, we denote by $\mathscr{C}^{0}(M,Y)$ the space of continuous maps $M\to Y$, and by $\mathscr{O}(M,Y)$ the space of holomorphic maps from some neighbourhood of $M$ in $X$ to $Y$. If $X$ is a Riemann surface, $Y\subset\mathbb{C}^{n}$ is a submanifold and $E\subset\mathring{M}$ is a finite subset, then
@@
denotes the free commutative group of finite divisors of the set $X$. (Note that we use multiplicative notation.) Given two divisors $D_{1},D_{2}\in\textrm{Div}(X)$, we write $D_{1}\geq D_{2}$ whenever the divisor $D_{1}D_{2}^{-1}$ is of the form $D_{1}D_{2}^{-1}=\prod_{i=1}^{\mathfrak{n}}p_{i}^{s_{i}}$ with $p_{i}\in X$, $s_{i}\geq 0$, $\mathfrak{n}\in\mathbb{N}$. In such case, we say that $D_{1}D_{2}^{-1}=D$ is an effective divisor. Moreover, the set $\mathrm{supp}(D)=\{p_{i}\colon s_{i}\neq 0\}$ is called the support of $D=\prod_{i=1}^{\mathfrak{n}}p_{i}^{s_{i}}\in\textrm{Div}(X)$.
@@
(See [6, Definition 1.12.9].) Given an open Riemann surface $M$, an admissible set in $M$ is a compact set $S=K\cup\Gamma$, where $K$ is a finite (or possibly empty) union of pairwise disjoint compact domains with piecewise smooth boundaries in $M$ and $\Gamma=\overline{S\setminus K}$ is a finite (or possibly empty) union of pairwise disjoint smooth Jordan arcs and closed Jordan curves intersecting $K$ only at their endpoints (or not at all) and such that their intersections with the boundary $bK$ of $K$ are transverse.
@@
    If $i,j\in\{1,\ldots,l\}$, $i\neq j$, and $C_{i}\cap C_{j}\neq\varnothing$, then $C_{i}\cap C_{j}=\{p_{i,j}\}$ where $p_{ij}$ is a common endpoint of $C_{i}$ and $C_{j}$.
@@
[…]
Let $M$ be an open Riemann surface, $S\subset M$ be a connected Runge admissible set, and $E\subset\mathring{S},\ \Lambda\subset S$ be disjoint finite subsets. Also let $A\subset\mathbb{C}^{n}$ be a closed irreducible conical complex subvariety and $\theta$ be a nowhere vanishing holomorphic $1$-form on $M$. Assume that $\mathscr{C}=\{C_{1},\dots,C_{l}\}$ is a skeleton of $S$ based at $E\cup\Lambda$ and set $C=\bigcup_{j=1}^{l}C_{j}$ (see Definition 2.5). For any map $f\in\mathscr{C}^{0}(C\setminus E,A_{*})$ we define the family of maps
@@
(2.5) $\mathscr{P}(h)$ ::=:= $\left(\int_{C_{j}}(h-f)\theta\right)_{j=1,\dots,l},\quad h\in\mathscr{C}^{0}(C,f).$
Recall that a complex manifold $X$ is called an Oka manifold if every holomorphic map $K\to X$ from a neighbourhood of a compact convex set $K\subset\mathbb{C}^{n}$ can be approximated uniformly on $K$ by entire maps $\mathbb{C}^{n}\to X$. (See [16, Definition 5.4.1].)
@@ 3.  The Mergelyan theorem for meromorphic $A$-immersions
In this section we establish Mergelyan and semiglobal approximation results for meromorphic $A$-immersions; see Theorem 3.1 and Proposition 3.2. The former, which is used to prove the latter under extra assumptions, can be seen as our main new technical tool.
@@
Let $M$ be an open Riemann surface, $S=K\cup\Gamma$ be an admissible set in $M$, and $E\subset\mathring{S}$, $\Lambda\subset S$ be disjoint finite subsets. Let $A\subset\mathbb{C}^{n}$ $(n\geq 3)$ be a closed irreducible conical subvariety which is not contained in any hyperplane and is smooth away from 00, and let $F\in\mathscr{I}_{A}(S|E)$. Then, for any $\varepsilon>0$ and any $r\in\mathbb{Z}_{+}$ there exist an open neighbourhood $\widetilde{S}\subset M$ of $S$ and a full meromorphic $A$-immersion $G\in\mathscr{I}_{A}(\widetilde{S}|E)$ such that the following are satisfied.
@@
    $G-F$ is continuous on $S$, vanishes on $\Lambda\setminus\mathring{S}$ and vanishes to order $r$ on $E\cup(\Lambda\cap\mathring{S})$.
@@
    The difference $G-F$ extends continuously to $S$, vanishes on $\Lambda\setminus\mathring{S}$ and vanishes to order $r$ on $E\cup(\Lambda\cap\mathring{S})$.
@@
    $||G-F||_{S}<\varepsilon$.
We begin with preparations for the proof of Theorem 3.1. Here is the key step.
@@
    $\tilde{f}-f$ is continuous on $S$ and $||\tilde{f}-f||_{S}<\varepsilon$.
@@
We proceed in two steps. Firstly, we approximate and interpolate $f$ by a full map $\hat{f}\in\mathscr{A}_{\infty}(S|E,A_{*})$ with $\mathscr{P}(\hat{f})=0$. Secondly, we find a period dominating spray of maps in $\mathscr{A}_{\infty}(S|E,A_{*})$ with core $\hat{f}$ and then approximate and interpolate it by a spray in $\widetilde{S}$. A suitable parameter value will then furnish a map $\tilde{f}\in\mathscr{O}_{\infty}(\widetilde{S}|E,A_{*})$ satisfying the conclusion of the lemma.
@@
Choose points $x_{1},\dots,x_{k}\in K\setminus(E\cup\Lambda)$ such that $T_{f(x_{j})}A$, $j=1,\dots,k$, span $\Sigma(f)$. Set $A^{\prime}:=\{0\}\cup\{z\in A_{*}\colon T_{z}A\subset\Sigma(f)\}$. Choose a holomorphic vector field $V$ on $A$ tangential to $A$ such that it vanishes at 00 and is not everywhere tangential to $A^{\prime}$ along $f(K\setminus E)$. Let $\mathbb{C}\ni t\mapsto\phi(t,z)\in A$ be the flow of $V$ for small values of $|t|$ and $\phi(0,z)=z\in A$. Fix a nonconstant function $h\in\mathscr{O}(K)$ with $h(x_{j})=0,\ j=1,\dots,k$, and vanishing to order $r$ at all points in $E\cup\Lambda$. For any function $g$ on a small neighbourhood of the constant zero-function in $\mathscr{A}(K)$, define $\Phi(g)\in\mathscr{A}_{\infty}\left(K|E,A_{*}\right)$ by
@@
where $\mathscr{P}$ is the period map in Lemma 3.3 (ii). We have $\mathscr{P}(\Phi(0))=\mathscr{P}(\phi(0,f))=\mathscr{P}(f)=0$ and since the vector space $\mathscr{A}(K)$ is of infinite dimension, there exists a nonconstant function $g\in\mathscr{A}(K)$ close to the zero-function on $K$, satisfying $\mathscr{P}(\Phi(g))=0$. It follows that the map $\hat{f}:=\Phi(g)\colon K\setminus E\to A_{*}$ lies in $\mathscr{A}_{\infty}(K|E,A_{*})$ and satisfies the requirements in Claim 3.4 except that $\hat{f}$ may fail to be full. Indeed, note first that if $p\in E\cup\Lambda$, then $(\hat{f}-f)(p)=\Phi(g)(p)-f(p)=\phi(0,f(p))-f(p)=f(p)-f(p)=0$ (to the correct order, as $h$ vanishes to order $r$ at $p$), and approximation on $K$ follows by construction.
Let us check that $\dim\Sigma(\hat{f})>\dim\Sigma(f)$, where $\Sigma(\hat{f})$ denotes the $\mathbb{C}$-linear subspace of $\mathbb{C}^{n}$ spanned by all tangent spaces $T_{z}A$ with $z\in\hat{f}(K\setminus E)$. Firstly, observe that $\Sigma(f)\subset\Sigma(\hat{f})$. Indeed, since $h(x_{j})=0,\ j=1,\dots,k$, the inclusion follows from the equality $\hat{f}(x_{j})=\Phi(g)(x_{j})=\phi(g(x_{j})h(x_{j}),f(x_{j}))=\phi(0,f(x_{j}))=f(x_{j})$. Moreover, $g$ is nonconstant on $K$ (hence so is $gh$) and $V$ is not everywhere tangential to $A^{\prime}$ along $f(K\setminus E)$, which implies that there exists a point $x_{0}\in K\setminus(E\cup\Lambda\cup\{x_{1},\dots,x_{k}\})$ such that $\hat{f}(x_{0})\in A\setminus A^{\prime}$. Therefore, $T_{\hat{f}(x_{0})}A\subset\Sigma(\hat{f})\setminus\Sigma(f)$ and $\dim\Sigma(\hat{f})>\dim\Sigma(f)$, as claimed. Finally, a finite recursive application of this argument enables us to assume that $\dim\Sigma(\hat{f})=n$, thereby concluding that $\hat{f}$ is full. ∎
@@
is continuous, holomorphic on $U\times(\mathring{S}\setminus E)$ and satisfies $\widetilde{\Phi}_{\tilde{f}_{0}}(0,\cdot)/g=\tilde{f}_{0}/g=f$. Moreover, $\widetilde{\Phi}_{\tilde{f}_{0}}(\zeta,\cdot)/g$ is meromorphic on $\mathring{S}$, the difference $\widetilde{\Phi}_{\tilde{f}_{0}}(\zeta,\cdot)/g-f$ extends continuously to $S$, hence $\widetilde{\Phi}_{\tilde{f}_{0}}(\zeta,\cdot)/g\in\mathscr{C}^{0}(C,f)$ (2.4), $\widetilde{\Phi}_{\tilde{f}_{0}}(\zeta,\cdot)/g-f$ vanishes to order $r$ on $E\cup(\Lambda\cap\mathring{S})$ for each parameter $\zeta\in U$, and the period map $U\ni\zeta\mapsto\mathscr{P}(\widetilde{\Phi}_{\tilde{f}_{0}}(\zeta,\cdot)/g)\in\mathbb{C}^{nl}$ (2) has maximal rank equal to $nl$ at $\zeta=0$.
@@
If the approximation of $f_{0}$ by $h_{0}$ is sufficiently close on $S$, then $\mathscr{R}(\zeta)=\mathscr{P}(\Phi_{H,h_{0}}(\zeta,\cdot)/g)\approx\mathscr{P}(\Phi_{H,\tilde{f}_{0}}(\zeta,\cdot)/g)=\widetilde{\mathscr{Q}}(\zeta)$ for each $\zeta\in U$, implying that $\mathscr{R}$ is submersive at $0\in\mathbb{C}^{Nl}$ and $0\in\mathbb{C}^{nl}$ lies in the range of $\mathscr{R}$ by the implicit function theorem. Thus, there exists a parameter $\zeta_{0}\in U$ close to 00 such that $\mathscr{R}(\zeta_{0})=0$. Defining $\tilde{f}:=\Phi_{H,h_{0}}(\zeta_{0},\cdot)/g\in\mathscr{O}_{\infty}(\widetilde{S}|E,A_{*})$, it satisfies the following properties.
@@
Assume that the admissible set $S$ is connected; if not, apply the same argument on each of its connected components. Moreover, assume that $K\neq\varnothing$; otherwise $S=\Gamma$ and $F\colon S\to\mathbb{C}^{n}$ is an $A$-immersion of class $\mathscr{C}^{1}(S)$ as in [4, Sect. 7], hence [4, Theorem 7.2 (a)] solves the problem except for the interpolation, which is achieved as in what follows. Fix a nowhere vanishing holomorphic $1$-form $\theta$ on $M$ and write $dF=f\theta$. Pick a point $p_{0}\in\mathring{S}\setminus(E\cup\Lambda)\subset\mathring{K}$ and let $\mathscr{C}=\{C_{1},\ldots,C_{l}\}$ be a skeleton of $S$ based at $E\cup\Lambda\cup\{p_{0}\}$ (see Definition 2.5). By Lemma 3.3, we may approximate $f$ uniformly on $S$ by a full meromorphic map $\tilde{f}\in\mathscr{A}_{\infty}(\widetilde{S}|E,A_{*})$ defined on a neighbourhood $\widetilde{S}$ of $S$ such that $S$ is a strong deformation retract of $\widetilde{S}$, $\tilde{f}$ agrees with $f$ to order $r$ on $E\cup(\Lambda\cap\mathring{S})$, and $\mathscr{P}(\tilde{f})=0$, where $\mathscr{P}$ is the period map (2) associated to $\mathscr{C}$, $f$ and $\theta$. In particular, $\tilde{f}\theta$ is exact on $\widetilde{S}$; indeed, $f\theta=dF$ and $(\tilde{f}-f)\theta$ are exact on $C=\bigcup_{i=1}^{l}C_{i}$, and $C$ is a strong deformation retract of $S$, hence of $\widetilde{S}$ (see the last part in the proof of Lemma 3.3). Therefore, the formula
@@
defines a full meromorphic $A$-immersion on $\widetilde{S}$ arbitrarily close to $F$ on $S$, and such that the difference $G-F$ is continuous on $S$, vanishes at all points in $\Lambda$ and vanishes to order $r+1$ on $E\cup(\Lambda\cap\mathring{S})$; take into account that $\mathscr{C}$ is based at $E\cup\Lambda\cup\{p_{0}\}$, $C$ is connected and $\mathscr{P}(\tilde{f})=0$. We conclude that $G$ satisfies the theorem. ∎

### v2

@@ 2. Notation and preliminaries
We write $\mathbb{N}=\{1,2,\ldots\}$ and $\mathbb{Z}_{+}=\mathbb{N}\cup\{0\}$. Let $|\cdot|$ denote the Euclidean norm in $\mathbb{R}^{n}$ for $n\in\mathbb{N}$. Given a compact topological space $K$ and a continuous map $f:K\to\mathbb{R}^{n}$, we mean by $\|f\|_{K}$ the maximum norm of $f$ on $K$. Given $f,g\colon K\to\mathbb{R}^{n}$, we say that $f$ approximates $g$ on $K$, and write $f\approx g$, if $f-g||_{K}<\varepsilon$ for some $\varepsilon>0$ which is sufficiently small according to the argument. Throughout the paper, we assume that surfaces are connected and $n\in\mathbb{N}$, $n\geq 3$.
Let $A\subset\mathbb{C}^{n}$ be a closed irreducible conical complex subvariety of $\mathbb{C}^{n}$ which is not contained in any hyperplane and is smooth away from 0.444A complex subvariety $A\subset\mathbb{C}^{n}$ is conical if $\lambda A=A$ for all $\lambda\in\mathbb{C}\setminus\{0\}$. We denote by $A_{*}:=A\setminus\{0\}$ the punctured subvariety. So, throughout the paper, we shall assume that $A_{*}$ is smooth and connected.
@@
Let $F=(F_{1},\dots,F_{n})\colon M\to\mathbb{C}^{n}$ be a meromorphic map which is holomorphic on $M\setminus E$ and has effective poles at all points in $E$. Fix a nowhere vanishing holomorphic $1$-form $\theta$ on $M$ (see [32]) and write $dF_{j}=f_{j}\theta$, $j=1,\dots,n$. Then the map $F$ is a meromorphic $A$-immersion $M\setminus E\to\mathbb{C}^{n}$ if and only if its derivative $f=dF/\theta=(dF_{1}/\theta,\dots,dF_{n}/\theta):M\setminus E\to\mathbb{C}^{n}$ has the range in $A_{*}$. On the other hand, given a meromorphic map $f=(f_{1},\dots,f_{n})\colon M\to A_{*}$ with effective poles in $E$, the vectorial $1$-form $f\theta=(f_{1}\theta,\dots,f_{n}\theta)$ integrates to a meromorphic $A$-immersion $F:M\setminus E\to\mathbb{C}^{n}$ in $\mathscr{I}_{A}(M|E)$ if and only if $f\theta$ is holomorphic and exact on $M\setminus E$. The corresponding meromorphic $A$-immersion is then obtained by
@@
Let $A$ be as in Definition 2.1, $M$ be either an open Riemann surface or a compact bordered Riemann surface, and $\theta$ be a nowhere vanishing holomorphic $1$-form on $M$. A holomorphic $A$-immersion $F\colon M\to\mathbb{C}^{n}$ is full if, setting $f=dF/\theta$, the $\mathbb{C}$-linear span of the tangent spaces $T_{f(x)}A,\ x\in M$, equals $\mathbb{C}^{n}$. Full meromorphic $A$-immersions are defined analogously.
Given complex manifolds X,YX,\ Y and a subset $M\subset X$, we denote by $\mathscr{C}^{0}(M,Y)$ the space of continuous maps $M\to Y$, and by $\mathscr{O}(M,Y)$ the space of holomorphic maps from some neighbourhood of $M$ in $X$ to $Y$. If $X$ is a Riemann surface, $Y\subset\mathbb{C}^{n}$ is a submanifold and $E\subset\mathring{M}$ is a finite subset, then
@@
the free commutative group of finite divisors of a set $X$. (Note that we use multiplicative notation.) Given two divisors $D_{1},D_{2}\in\textrm{Div}(X)$, we write $D_{1}\geq D_{2}$ whenever the divisor $D_{1}D_{2}^{-1}$ is of the form $D_{1}D_{2}^{-1}=\prod_{i=1}^{\mathfrak{n}}p_{i}^{s_{i}}$ with $p_{i}\in X$, $s_{i}\geq 0$, $\mathfrak{n}\in\mathbb{N}$. In such case, we say that $D_{1}D_{2}^{-1}=D$ is an effective divisor. Moreover, the set $\mathrm{supp}(D)=\{p_{i}\colon s_{i}\neq 0\}$ is called the support of $D=\prod_{i=1}^{\mathfrak{n}}p_{i}^{s_{i}}\in\textrm{Div}(X)$.
@@
(See [7, Definition 1.12.9].) Given an open Riemann surface $M$, an admissible set in $M$ is a compact set $S=K\cup\Gamma$, where $K$ is a finite (or possibly empty) union of pairwise disjoint compact domains with piecewise smooth boundaries in $M$ and $\Gamma=\overline{S\setminus K}$ is a finite (or possibly empty) union of pairwise disjoint smooth Jordan arcs and closed Jordan curves intersecting $K$ only at their endpoints (or not at all) and such that their intersections with the boundary $bK$ of $K$ are transverse.
@@
    If $\in\{1,\ldots,l\}, iji\neq j$, and $C_{i}\cap C_{j}\neq\varnothing$, then $C_{i}\cap C_{j}=\{p_{i,j}\}$ where $p_{ij}$ is a common endpoint of $C_{i}$ and $C_{j}$.
@@
Given a connected admissible set $S$, one can always find skeletons of $S$ based at any finite set $Q\subset S$ as in the definition. For a detailed construction of a family of such curves we refer to the proof of [7, Proposition 3.3.2]. We emphasize that by property $\rm(S1)$, $C$ is Runge in $M$ whenever $S$ is so. Recall that a compact set $K$ in a Riemann surface $M$ is holomorphically convex or a Runge set in $M$ if and only if $K$ has no holes in $M$ (see [7, Lemma 1.12.3]).
Let $M$ be an open Riemann surface, $S\subset M$ be a connected Runge admissible set (that is, $S$ is an admissible set which is in addition a Runge subset of $M$), and $E\subset\mathring{S},\ \Lambda\subset S$ be disjoint finite subsets. Also let $A\subset\mathbb{C}^{n}$ be a closed irreducible conical complex subvariety and $\theta$ be a nowhere vanishing holomorphic $1$-form on $M$. Assume that $\mathscr{C}=\{C_{1},\dots,C_{l}\}$ is a skeleton of $S$ based at $E\cup\Lambda$ and set $C:=\bigcup_{j=1}^{l}C_{j}$ (see Definition 2.5). For any map $f\in\mathscr{C}^{0}(C\setminus E,A_{*})$ we define the family of maps
@@
(2.5) $\mathscr{P}(h)$ $:=$ $\left(\int_{C_{j}}(h-f)\theta\right)_{j=1,\dots,l},\quad h\in\mathscr{C}^{0}(C,f).$
As already mentioned in the introduction, a special class of complex manifolds, called Oka manifolds, plays an important role in the results of the present paper. They were firstly introduced by Forstnerič in [23], following the works by Oka, Grauert, and Gromov on the classical Oka-Grauert principle [31]. We say that a complex manifold $X$ is an Oka manifold if every holomorphic map $K\to X$ from (a neighbourhood of) any compact convex set $K\subset\mathbb{C}^{n}$ for any $n\in\mathbb{N}$ can be approximated uniformly on $K$ by entire maps $\mathbb{C}^{n}\to X$ [24, Definition 5.4.1]. The main result in this theory is that maps from a Stein manifold to an Oka manifold satisfy all equivalent forms of the Oka principle [24, Theorem 5.4.4]. This includes the Oka property with approximation and interpolation to the effect that given a compact Runge set $K$ in a Stein manifold $M$ and an Oka manifold $X$, every continuous map $f:M\to X$ which is holomorphic on a neighbourhood of $K$ can be approximated uniformly on $K$, and interpolated to any given finite order on a finite set in $K$, by holomorphic maps $M\to X$ that are homotopic to $f$. This shall be a crucial tool in our method of proof.
@@ 3. The Mergelyan theorem for meromorphic $A$-immersions
In this section we establish Mergelyan and semiglobal approximation results for meromorphic $A$-immersions; see Theorem 3.1 and Proposition 3.2. The former, which is used to prove the latter under extra assumptions, can be seen as our main new technical tool. Oka theory is not required for this task, so the submanifold $A_{*}$ in Theorem 3.1 need not be Oka. On the other hand, we provide a similar result of semiglobal nature in Proposition 3.2, under the extra assumption that $A_{*}$ is Oka.
@@
Let $M$ be an open Riemann surface, $S=K\cup\Gamma$ be an admissible set in $M$, and $E\subset\mathring{S}$, $\Lambda\subset S$ be disjoint finite subsets. Let $A\subset\mathbb{C}^{n}$ $(n\geq 3)$ be a closed irreducible conical subvariety which is not contained in any hyperplane and is smooth away from 0, and let $F\in\mathscr{I}_{A}(S|E)$. Then, for any $\varepsilon>0$ and any $r\in\mathbb{Z}_{+}$ there exist an open neighbourhood $\widetilde{S}\subset M$ of $S$ and a full meromorphic $A$-immersion $G\in\mathscr{I}_{A}(\widetilde{S}|E)$ such that the following are satisfied.
@@
    GFG-F is continuous on $S$, vanishes on $\Lambda\setminus\mathring{S}$ and vanishes to order $r$ on $E\cup(\Lambda\cap\mathring{S})$.
@@
    The difference GFG-F extends continuously to $S$, vanishes on $\Lambda\setminus\mathring{S}$ and vanishes to order $r$ on $E\cup(\Lambda\cap\mathring{S})$.
@@
    $G-F||_{S}<\varepsilon$.
We begin with preparations for the proof of Theorem 3.1. The following lemma is the key step; a Mergelyan type theorem with interpolation for meromorphic maps into a conical subvariety $A_{*}\subset\mathbb{C}^{n}$. Its main feature is the control of the periods of the approximating meromorphic map, which ensures that it indeed integrates to a directed meromorphic immersion.
@@
    $\tilde{f}-f$ is continuous on $S$ and $\tilde{f}-f||_{S}<\varepsilon$.
@@
The proof consists of adapting the method of sprays developed in [4] to the meromorphic framework. We proceed in two steps. Firstly, we approximate and interpolate $f$ by a full map $\hat{f}\in\mathscr{A}_{\infty}(S|E,A_{*})$ with $\mathscr{P}(\hat{f})=0$. Secondly, we find a period dominating spray of maps in $\mathscr{A}_{\infty}(S|E,A_{*})$ with core $\hat{f}$ and then approximate and interpolate it by a spray in $\widetilde{S}$. A suitable parameter value will then furnish a map $\tilde{f}\in\mathscr{O}_{\infty}(\widetilde{S}|E,A_{*})$ satisfying the conclusion of the lemma.
@@
Choose points $x_{1},\dots,x_{k}\in K\setminus(E\cup\Lambda)$ such that $T_{f(x_{j})}A$, $j=1,\dots,k$, span $\Sigma(f)$. Set $A^{\prime}:=\{0\}\cup\{z\in A_{*}\colon T_{z}A\subset\Sigma(f)\}$. Choose a holomorphic vector field $V$ on $A$ tangential to $A$ such that it vanishes at 0 and is not everywhere tangential to $A^{\prime}$ along $f(K\setminus E)$. Let $\mathbb{C}\ni t\mapsto\phi(t,z)\in A$ be the flow of $V$ for small values of $|t|$ and $\phi(0,z)=z\in A$. Fix a nonconstant function $h\in\mathscr{O}(K)$ with $h(x_{j})=0,\ j=1,\dots,k$, and vanishing to order $r$ at all points in $E\cup\Lambda$. For any function $g$ on a small neighbourhood of the constant zero-function in $\mathscr{A}(K)$, define $\Phi(g)\in\mathscr{A}_{\infty}\left(K|E,A_{*}\right)$ by
@@
where $\mathscr{P}$ is the period map in Lemma 3.3-(ii). We have $\mathscr{P}(\Phi(0))=\mathscr{P}(\phi(0,f))=\mathscr{P}(f)=0$ and since the vector space $\mathscr{A}(K)$ is of infinite dimension, there exists a nonconstant function $g\in\mathscr{A}(K)$ close to the zero-function on $K$, satisfying $\mathscr{P}(\Phi(g))=0$. It follows that the map $\hat{f}:=\Phi(g)\colon K\setminus E\to A_{*}$ lies in $\mathscr{A}_{\infty}(K|E,A_{*})$ and satisfies the requirements in Claim 3.4 except that $\hat{f}$ may fail to be full. Indeed, note first that if $p\in E\cup\Lambda$, then $(\hat{f}-f)(p)=\Phi(g)(p)-f(p)=\phi(0,f(p))-f(p)=f(p)-f(p)=0$ (to the correct order, as $h$ vanishes to order $r$ at $p$), and approximation on $K$ follows by construction.
[…]
@@
is continuous, holomorphic on $U\times(\mathring{S}\setminus E)$ and satisfies $\widetilde{\Phi}_{\tilde{f}_{0}}(0,\cdot)/g=\tilde{f}_{0}/g=f$. Moreover, $\widetilde{\Phi}_{\tilde{f}_{0}}(\zeta,\cdot)/g$ is meromorphic on $\mathring{S}$, the difference $\widetilde{\Phi}_{\tilde{f}_{0}}(\zeta,\cdot)/g-f$ extends continuously to $S$, hence $\widetilde{\Phi}_{\tilde{f}_{0}}(\zeta,\cdot)/g\in\mathscr{C}^{0}(C,f)$ (2.4), $\widetilde{\Phi}_{\tilde{f}_{0}}(\zeta,\cdot)/g-f$ vanishes to order $r$ on $E\cup(\Lambda\cap\mathring{S})$ for each parameter $\zeta\in U$, and the period map $U\ni\zeta\mapsto\mathscr{P}(\widetilde{\Phi}_{\tilde{f}_{0}}(\zeta,\cdot)/g)\in\mathbb{C}^{nl}$ (2) has maximal rank equal to nlnl at $\zeta=0$.
@@
If the approximation of $f_{0}$ by $h_{0}$ is sufficiently close on $S$, then $\mathscr{R}(\zeta)=\mathscr{P}(\Phi_{H,h_{0}}(\zeta,\cdot)/g)\approx\mathscr{P}(\Phi_{H,\tilde{f}_{0}}(\zeta,\cdot)/g)=\widetilde{\mathscr{Q}}(\zeta)$ for each $\zeta\in U$, implying that $\mathscr{R}$ is submersive at $0\in\mathbb{C}^{Nl}$ and $0\in\mathbb{C}^{nl}$ lies in the range of $\mathscr{R}$ by the implicit function theorem. Thus, there exists a parameter $\zeta_{0}\in U$ close to 0 such that $\mathscr{R}(\zeta_{0})=0$. Defining $\tilde{f}:=\Phi_{H,h_{0}}(\zeta_{0},\cdot)/g\in\mathscr{O}_{\infty}(\widetilde{S}|E,A_{*})$, it satisfies the following properties.
@@
The idea of the proof is just to pass to the derivative of the original directed meromorphic immersion, use Lemma 3.3, and then integrate the map furnished by the lemma to obtain a directed meromorphic immersion meeting conclusions of the theorem.
Assume that the admissible set $S$ is connected; if not, apply the same argument on each of its connected components. Moreover, assume that $K\neq\varnothing$; otherwise $S=\Gamma$ and $F\colon S\to\mathbb{C}^{n}$ is an $A$-immersion of class $\mathscr{C}^{1}(S)$ as in [4, §7], hence [4, Theorem 7.2-(a)] solves the problem except for the interpolation, which is achieved as in what follows. Fix a nowhere vanishing holomorphic $1$-form $\theta$ on $M$ and write $dF=f\theta$. Pick a point $p_{0}\in\mathring{S}\setminus(E\cup\Lambda)\subset\mathring{K}$ and let $\mathscr{C}=\{C_{1},\ldots,C_{l}\}$ be a skeleton of $S$ based at $E\cup\Lambda\cup\{p_{0}\}$ (see Definition 2.5). By Lemma 3.3, we may approximate $f$ uniformly on $S$ by a full meromorphic map $\tilde{f}\in\mathscr{A}_{\infty}(\widetilde{S}|E,A_{*})$ defined on a neighbourhood $\widetilde{S}$ of $S$ such that $S$ is a strong deformation retract of $\widetilde{S}$, $\tilde{f}$ agrees with $f$ to order $r$ on $E\cup(\Lambda\cap\mathring{S})$, and $\mathscr{P}(\tilde{f})=0$, where $\mathscr{P}$ is the period map (2) associated to $\mathscr{C}$, $f$ and $\theta$. In particular, $\tilde{f}\theta$ is exact on $\widetilde{S}$; indeed, $f\theta=dF$ and $(\tilde{f}-f)\theta$ are exact on $C=\bigcup_{i=1}^{l}C_{i}$, and $C$ is a strong deformation retract of $S$, hence of $\widetilde{S}$ (see the last part in the proof of Lemma 3.3). Therefore, the formula
@@
defines a full meromorphic $A$-immersion on $\widetilde{S}$ arbitrarily close to $F$ on $S$, and such that the difference GFG-F is continuous on $S$, vanishes at all points in $\Lambda$ and vanishes to order $r+1$ on $E\cup(\Lambda\cap\mathring{S})$; take into account that $\mathscr{C}$ is based at $E\cup\Lambda\cup\{p_{0}\}$, $C$ is connected and $\mathscr{P}(\tilde{f})=0$. We conclude that $G$ satisfies the theorem. ∎